# library.bib

### v1

@@ Appendix C Models
Kaplan-Meier (KM) [kaplanmeier1958] kaplan survival v3.5.7 ExpMort
Nelson-Aalen (NA) [aalen1978] nelson survival v3.5.7 ExpMort
Akritas Estimator (AK) [akritas1994] akritas survivalmodels v0.1.18 ExpMort
Cox PH (CPH) [cox1975] coxph survival v4.1.8 ✓ (Breslow) lp
CV Regularized CPH (GLMN) [simon2011] cv_glmnet glmnet v4.1-8 ✓ (Breslow) lp
Penalized (Pen) [goeman2010l1penalized] penalized penalized v0.9.52 ✓ (Breslow) ExpMort
Parametric (AFT)  [kalbfleisch2011] parametric survival v3.5.7 ✓ (AFT) lp
Flexible Splines (Flex) [roystonparmar2002] flexible flexsurv v2.2.2 lp
Random Survival Forest (RFSRC) [ishwaran2008] rfsrc randomForestSRC v3.2.2 ExpMort
Random Survival Forest (RAN) [ishwaran2008, pkgranger] ranger ranger v0.16.0 ExpMort
Conditional Inference Forest (CIF) [hothorn2006] cforest partykit v1.2.20 ExpMort
Oblique Random Survival Forest (ORSF) [jaeger2024] orsf aorsf v0.1.2 ExpMort
Relative Risk Tree (RRT) [breiman1984] rpart rpart v4.1.23 (KM, PH)
Model-Based Boosting (MBST) [buhlmann2003boostingl2] mboost mboost v.2.9.9 ✓ (Breslow) lp
CoxBoost (CoxB) [binder2008allowing] cv_coxboost CoxBoost v1.5 ✓ (Breslow) lp
XGBoost (XGBCox) [chen2016xgboost] xgboost xgboost v1.7.6.1 ✓ (Breslow) lp
XGBoost (XGBAFT) [barnwal2022] xgboost xgboost v1.7.6.1 (KM, AFT) lp
SSVM-Hybrid (SSVM) [vanbelle2011supportvector] svm survivalsvm v0.0.5 (KM, PH)
@@
  • 2. Learner ID in mlr3. Most learners are implemented in mlr3extralearners v0.7.1-9000 with KM, NA, CPH residing in mlr3proba v0.6.0
  • 3. Package in which the learner is implemented with version used in experiment.
  • 4. distr predict type in mlr3proba is the probabilistic prediction. A check (✓) represents the distribution being predicted directly by the package, in some cases using the Breslow estimator. The KM estimator may otherwise be used to estimate a baseline distribution from which the predicted distribution is composed with either PH or AFT forms, unless the Breslow estimator is used directly to obtain distribution predictions from the linear predictors. Notably for MBST, the availability of distr depends on the family parameter.
  • 5. crank predict type in mlr3proba is the continuous ranking prediction. A check (✓) represents the ranking being predicted directly by the package. ‘ExpMort’ stands for expected mortality and is a risk score composed from the predicted survival distribution (distr). ‘lp’ represents the ranking being identical to the predicted linear predictor.
@@
Appendix D Model Configuration
LABEL:tab:config shows hyperparameter search spaces and non-default parameter values as well as common pre-processing requirements.
@@ Appendix D Model Configuration
MBST family mstop nu baselearner {gehan, cindex, coxph,weibull} $\{10,...,5000\}$ $(0,0.1]$bols, btree}
@@
XGBCox objective tree_method booster max_depth subsample colsample_bytree nrounds eta grow_policy survival:cox hist gbtree {1,…,20} $[0,1]$ $(0,1]$10,…,5000} $[10^{-5},10^{5}]$depthwise, lossguide}
XGBAFT objective tree_method booster max_depth subsample colsample_bytree nrounds eta grow_policy aft_loss_distribution aft_loss_distribution_scale survival:aft hist gbtree {1,…,20} $[0,1]$ $(0,1]$10,…,5000} $[10^{-5},10^{5}]$depthwise, lossguide} {normal, logistic, extreme} $[0.5,2]$
SSVM5 type diff.meth gamma.mu kernel kernel.pars hybrid makediff3 $([2^{-10},2^{10}],[2^{-10},2^{10}])$lin_kernel, rbf_kernel, add_kernel} $[2^{-5},2^{5}]$
@@
Appendix E Implementation, Reproducibility, Accessibility
Platform All experiments were conducted on R 4.2.2 (2022-10-31) – “Innocent and Trusting” on the Beartooth Computing Environment [beartooth].
Reproducibility and Accessibility Seeds were set with L’Ecuyer’s random number generator [lecuyer1999] to ensure reproducible results. All code required to run the experiments, as well as the results, are freely available in a public GitHub repository (https://github.com/slds-lmu/paper_2023_survival_benchmark. Software, packages, and version numbers that were utilized to conduct benchmark and analysis are listed below. We also employ the renv R package [pkgrenv] to record and restore package dependencies to further aid reproducibility.
Packages All code is implemented using R v4.2.2 [pkgr] and the experiment was run with batchtools [lang2017batchtoolstools]. Models are implemented in mlr3proba v0.6.0 [pkgmlr3proba], mlr3extralearners v0.7.1-9000 [pkgmlr3extralearners] and survivalmodels v0.1.18 [pkgsurvivalmodels]. Tuning is implemented in mlr3tuning v0.19.2 [pkgmlr3tuning]. Benchmarking functionality is implemented in mlr3 v0.17.2 [pkgmlr3]. Benchmark analysis is implemented in mlr3benchmark v0.1.6 [pkgmlr3benchmark]. Compositions and pre-processing steps are implemented through mlr3pipelines v0.5.0-2 [pkgmlr3pipelines]. The packages and versions for the tested models are given in LABEL:tab:models, all measures are implemented in mlr3proba.
Appendix F Results
The following plots show boxplots of the respective evaluation measure across the five outer resampling folds. Results are also available following links on the GitHub repository at https://github.com/slds-lmu/paper_2023_survival_benchmark.
As described in Section 4, we provide aggregated results based on average ranks across all datasets as boxplots with two versions for discrimination measures and three different scaling options for scoring rules:
@@ Appendix F Results
F.1 Discrimination Measures
F.1.1 Raw Scores
(a) Models tuned and evaluated on Harrell’s
(b) Models tuned for Harrell’s C and evaluated on Uno’s C
(c) Models tuned for Harrell’s C and evaluated on ISBS
@@ F.1.1 Raw Scores
Figure 3: Boxplots of raw evaluation scores using discrimination measures for tuning (Harrell’s C) and using discrimination measures and ISBS for evaluation.
F.1.2 Scaled
(a) Models tuned and evaluated on Harrell’s
(b) Models tuned for Harrell’s C and evaluated on Uno’s C
(c) Models tuned for Harrell’s C and evaluated on ISBS
Figure 4: Boxplots of scaled evaluation scores using discrimination measures for tuning (Harrell’s C) and using discrimination measures and ISBS for evaluation.
F.2 Scoring Rules
F.2.1 Raw Scores
(a) Scores for models tuned and evaluated on RCLL
(b) Scores for models tuned for RCLL and evaluated on RISLL
(c) Scores for models tuned for RCLL and evaluated on RNLL
@@ F.2.1 Raw Scores
(d) Scores for models tuned for RCLL and evaluated on ISBS
Figure 5: Boxplots of raw evaluation scores using scoring rules for tuning (RCLL) and evaluation.
F.2.2 ERV
(a) ERV scores for models tuned and evaluated on RCLL
(b) ERV scores for models tuned for RCLL and evaluated on RISLL
(c) ERV scores for models tuned for RCLL and evaluated on RNLL
(d) ERV scores for models tuned for RCLL and evaluated on ISBS
Figure 6: Boxplots of ERV evaluation scores using scoring rules for tuning (RCLL) and evaluation.
F.3 Errors
@@
During the many computational steps performed during this benchmark, software errors are inevitably bound to happen. As we noted in Section 3, we impute missing evaluation scores in resampling folds using the score of the Kaplan-Meier estimator, which affects the results presented in this paper.
@@
Table 6 counts the number of errors encountered for each model, dataset, and tuning measures uses in the benchmark per outer resampling fold (up to five). We note that particularly the tasks hdfail and child have caused the majority of the runtime- or memory-related errors here due to their large sample sizes and number of unique time points.
@@
Table 6: Number of errors per outer resampling iteration (up to five), separated by model, dataset, and tuning measure.
Model Dataset Harrell’s C RCLL Total Errors
AK child 4 4 8
AK hdfail 5 5 10
GLMN hdfail 0 1 1
Pen child 0 2 2
Pen hdfail 4 0 4
RFSRC child 0 1 1
RFSRC hdfail 1 1 2
RFSRC nafld1 4 5 9
RAN child 4 5 9
RAN colrec 3 5 8
RAN flchain 0 1 1
RAN hdfail 4 5 9
RAN metabric 2 2 4
RAN nafld1 4 5 9
RAN support 4 5 9
CIF colrec 1 0 1
CIF hdfail 2 2 4
CIF nafld1 5 5 10
CIF support 2 2 4
ORSF nafld1 1 5 6
RRT hdfail 0 2 2
MBST aids2 0 3 3
MBST child 5 5 10
MBST dataFTR 1 1 2
MBST flchain 3 2 5
MBST gbsg 1 0 1
MBST hdfail 5 5 10
MBST metabric 1 0 1
MBST nafld1 5 5 10
MBST nwtco 4 5 9
MBST support 4 1 5
CoxB hdfail 5 5 10
XGBAFT hdfail 0 1 1
Appendix G Results per Dataset
@@
G.1 Discrimination
Figure 7: Raw scores for Harrell’s C across outer resampling folds per dataset, task, and evaluation measure for models tuned on Harrell’s C
Figure 8: Raw scores for Uno’s C across outer resampling folds per dataset, task, and evaluation measure for models tuned on Harrell’s C
Figure 9: Raw scores for ISBS across outer resampling folds per dataset, task, and evaluation measure for models tuned on Harrell’s C
G.2 Scoring Rules
Figure 10: Raw scores for RCLL across outer resampling folds per dataset, task, and evaluation measure for models tuned on RCLL
Figure 11: Raw scores for ISBS across outer resampling folds per dataset, task, and evaluation measure for models tuned on RCLL
Figure 12: Raw scores for RNLL across outer resampling folds per dataset, task, and evaluation measure for models tuned on RCLL
Figure 13: Raw scores for RISLL across outer resampling folds per dataset, task, and evaluation measure for models tuned on RCLL
G.3 Calibration
Both calibration measures here are presented only for models tuned on RCLL. We note that these measures may provide additional insights, yet consider both to be experimental in theory and implementation.
G.3.1 D-Calibration
D-Calibration considers a model well-calibrated if the p-value of the underlying test is greater than 0.05, which we display in form of a heatmap with X indicating a significant test result, indicating a model is not well-calibrated.
Figure 14: D-Calibration p-value heatmaps across all datasets and learners. ‘X’ indicates $p<0.05$ while a non-significant result indicates good calibration.
G.3.2 Alpha-Calibration
Van Houwelingen’s $\alpha$ relates predicted and observed hazards, with a value close to 1 implying a well-calibrated model.
Figure 15: Calibration scores using van Houwelingen’s $\alpha$ indicate good calibration when values are close to 1.

### v2

@@ Appendix C Models and Configurations
Kaplan-Meier (KM) (Kaplan and Meier, 1958) kaplan survival ExpMort
Nelson-Aalen (NEL) (Aalen, 1978) nelson survival ExpMort
Akritas Estimator (AK) (Akritas, 1994) akritas survivalmodels ExpMort
Cox PH (CPH) (Cox, 1972) coxph survival ✓ (Breslow) lp
CV Regularized CPH (GLMN) (Simon et al., 2011) cv_glmnet glmnet ✓ (Breslow) lp
Penalized (Pen) (Goeman, 2010) penalized penalized ✓ (Breslow) ExpMort
Parametric (AFT)  (Kalbfleisch and Prentice, 2011) parametric survival AFT lp
Flexible Splines (Flex) (Royston and Parmar, 2002) flexible flexsurv lp
Random Survival Forest (RFSRC) (Ishwaran et al., 2008) rfsrc randomForestSRC ExpMort
Random Survival Forest (RAN) (Ishwaran et al., 2008; Wright and Ziegler, 2017) ranger ranger ExpMort
Conditional Inference Forest (CIF) (Hothorn et al., 2006) cforest partykit ExpMort
Oblique Random Survival Forest (ORSF) (Jaeger et al., 2024) orsf aorsf ExpMort
Relative Risk Tree (RRT) (Breiman, 1998) rpart rpart
Model-Based Boosting (MBSTCox) (Bühlmann and Yu, 2003) mboost mboost ✓ (Breslow) lp
Model-Based Boosting (MBSTAFT) (Bühlmann and Yu, 2003) mboost mboost lp
CoxBoost (CoxB) (Binder and Schumacher, 2008) cv_coxboost CoxBoost ✓ (Breslow) lp
XGBoost (XGBCox) (Chen and Guestrin, 2016) xgboost xgboost Breslow lp
XGBoost (XGBAFT) (Barnwal et al., 2022) xgboost xgboost lp
SSVM-Hybrid (SSVM) (Van Belle et al., 2011) svm survivalsvm
@@
  • 2. Learner ID in mlr3.
  • 3. Package in which the learner is implemented. Most learners are available in mlr3extralearners, with KM, RRT, and CPH provided via mlr3proba, where existing R implementations are wrapped into the mlr3 framework. AK is natively implemented in survivalmodels.
  • 4. distr predict type in mlr3proba is the probabilistic prediction. A check (✓) incidcates that the distribution is provided directly by the package, somtimes via the Breslow estimator (in parentheses). Parametric (AFT) models construct the distribution internally in survivalmodels, while XGBCox relies on mlr3proba ’s breslow() function since xgboost does not estimate distributions by default. A cross (✕) means the prediction is not available.
  • 5. crank predict type in mlr3proba is the continuous ranking prediction. A check (✓) indicates that the ranking is being predicted directly by the package. ‘ExpMort’ stands for expected mortality and is a risk score composed from the predicted survival distribution (distr). ‘lp’ represents the ranking being identical to the predicted linear predictor.
@@
MBSTCox family mstop nu baselearner coxph $\{10,...,5000\}$ $(0,0.1]$ {bols, btree}
MBSTAFT family mstop nu baselearner {gehan, weibull} $\{10,...,5000\}$ $(0,0.1]$ {bols, btree}
@@
XGBCox objective tree_method booster max_depth subsample colsample_bytree early_stopping_rounds eta grow_policy survival:cox hist gbtree {1,…,20} $[0,1]$ $(0,1]$ 50 $10^{-5},10^{5}]$ {depthwise, lossguide}
XGBAFT objective tree_method booster max_depth subsample colsample_bytree early_stopping_rounds eta grow_policy aft_loss_distribution aft_loss_distribution_scale survival:aft hist gbtree {1,…,20} $[0,1]$ $(0,1]$ 50 $10^{-5},10^{5}]$ {depthwise, lossguide} {normal, logistic, extreme} $[0.5,2]$
SSVM5 type diff.meth gamma.mu kernel kernel.pars hybrid makediff3 $([2^{-10},2^{10}],[2^{-10},2^{10}])$ {lin_kernel, rbf_kernel, add_kernel} $2^{-5},2^{5}]$
@@
Appendix D Implementation, Reproducibility, Accessibility
Platform All experiments were conducted on R 4.4.3 (2025-02-28) – “Trophy Case” on the Beartooth Computing Environment (Beartooth Computing Environment, x86_64 cluster, 2024).
Reproducibility and Accessibility Seeds were set with L’Ecuyer’s random number generator (L’Ecuyer, 1999) to ensure reproducible results. All code required to run the experiments, as well as the results, are freely available in a public GitHub repository (https://github.com/slds-lmu/paper_2023_survival_benchmark. Software, packages, and version numbers that were utilized to conduct benchmark and analysis are listed below. We also employ the renv R package (Ushey and Wickham, 2024) to record and restore package dependencies to further aid reproducibility.
Packages All code is implemented in R and the experiment was run with batchtools (Lang et al., 2017). Learners are implemented in mlr3proba (Sonabend et al., 2021), mlr3extralearners (Sonabend and Schratz, 2024) and survivalmodels (Sonabend, 2020). Tuning is implemented in mlr3tuning (Lang et al., 2023). Benchmarking functionality is implemented in mlr3 (Lang et al., 2019). Statistical benchmark analysis is implemented in mlr3benchmark (Sonabend and Pfisterer, 2020). Compositions and pre-processing steps are implemented through mlr3pipelines (Binder et al., 2021). The implementing packages for all algorithms are given in LABEL:tab:models and all measures are implemented in mlr3proba.
Appendix E Results
The following figures show boxplots of the respective evaluation measure across all outer resampling iterations. Results are also available following links on the GitHub repository at https://github.com/slds-lmu/paper_2023_survival_benchmark.
As described in Section˜4, we provide aggregated results based on average ranks across all datasets as boxplots with two versions for discrimination measures and three different scaling options for scoring rules:
@@ Appendix E Results
E.1 Discrimination Measures
E.1.1 Raw Scores
@@
Boxplots of raw evaluation scores using discrimination measures for tuning (Harrell’s C)
Figure 5: Learners tuned and evaluated with Harrell’s C
Figure 6: Learners tuned for Harrell’s C and evaluated with Uno’s C
E.1.2 Scaled
Figure 7: Learners tuned and evaluated with Harrell’s C (scaled scores)
Figure 8: Learners tuned for Harrell’s C and evaluated with Uno’s C (scaled scores)
E.2 Scoring Rules
E.2.1 Raw Scores
Figure 9: Learners tuned for ISBS and evaluated with ISBS (raw scores)
@@ E.2.1 Raw Scores
Figure 10: Learners tuned for ISBS and evaluated with ISLL (raw scores)
E.2.2 ERV
Figure 11: Learners tuned for ISBS and evaluated with ISBS (ERV)
Figure 12: Learners tuned for ISBS and evaluated with ISLL (ERV)
E.2.3 Scaled
Figure 13: Learners tuned for ISBS and evaluated with ISBS (scaleds)
Figure 14: Learners tuned for ISBS and evaluated with ISLL (scaled)
E.3 Errors
@@
During the many computational steps performed during this benchmark, software errors are inevitably bound to happen. As we noted in Section˜3, we impute missing evaluation scores in resampling folds using the score of the Kaplan-Meier estimator, which affects the results presented in this paper.
@@
Table LABEL:tab:errors counts the number of errors encountered for each model, dataset, and tuning measures uses in the benchmark per outer resampling fold (up to five). We note that particularly the tasks hdfail and child have caused the majority of the runtime- or memory-related errors here due to their large sample sizes and number of unique time points.
@@
Table 6: Number of errors per outer resampling iteration (up to 30), separated by model, dataset, and tuning measure.
Model Dataset Harrell’s C ISBS Total Errors
AK CarpenterFdaData 0 (0%) 1 (3.3%) 1
AK channing 1 (3.3%) 1 (3.3%) 2
AK child 3 (100%) 3 (100%) 6
AK e1684 3 (10%) 0 (0%) 3
AK hdfail 3 (100%) 3 (100%) 6
AK lung 8 (26.7%) 0 (0%) 8
AK uis 2 (6.7%) 0 (0%) 2
AK veteran 3 (10%) 0 (0%) 3
GLMN bladder0 1 (3.3%) 0 (0%) 1
GLMN channing 0 (0%) 1 (3.3%) 1
GLMN check_times 2 (66.7%) 0 (0%) 2
GLMN cost 12 (40%) 0 (0%) 12
GLMN dataSTR 2 (6.7%) 0 (0%) 2
GLMN hdfail 0 (0%) 3 (100%) 3
GLMN std 6 (20%) 0 (0%) 6
GLMN uis 4 (13.3%) 0 (0%) 4
GLMN veteran 0 (0%) 14 (46.7%) 14
GLMN wbc1 0 (0%) 4 (13.3%) 4
Pen aids.id 1 (3.3%) 9 (30%) 10
Pen bladder0 8 (26.7%) 0 (0%) 8
Pen channing 1 (3.3%) 0 (0%) 1
Pen check_times 3 (100%) 3 (100%) 6
Pen cost 3 (10%) 0 (0%) 3
Pen dataSTR 11 (36.7%) 3 (10%) 14
Pen hdfail 0 (0%) 2 (66.7%) 2
Flex aids.id 0 (0%) 10 (33.3%) 10
Flex check_times 3 (100%) 3 (100%) 6
Flex child 3 (100%) 3 (100%) 6
Flex dataFTR 2 (6.7%) 0 (0%) 2
Flex hdfail 3 (100%) 3 (100%) 6
Flex lung 9 (30%) 0 (0%) 9
Flex nafld1 14 (93.3%) 14 (93.3%) 28
Flex nwtco 15 (100%) 15 (100%) 30
Flex support 3 (100%) 3 (100%) 6
Flex wa_churn 15 (100%) 15 (100%) 30
RFSRC check_times 3 (100%) 3 (100%) 6
RFSRC child 3 (100%) 3 (100%) 6
RFSRC colrec 1 (33.3%) 2 (66.7%) 3
RFSRC nafld1 2 (13.3%) 1 (6.7%) 3
RFSRC support 2 (66.7%) 3 (100%) 5
RAN check_times 3 (100%) 2 (66.7%) 5
RAN child 3 (100%) 3 (100%) 6
RAN cost 1 (3.3%) 0 (0%) 1
RAN hdfail 1 (33.3%) 1 (33.3%) 2
RAN mgus 0 (0%) 2 (6.7%) 2
RAN nafld1 4 (26.7%) 9 (60%) 13
CIF child 3 (100%) 3 (100%) 6
CIF hdfail 3 (100%) 3 (100%) 6
ORSF child 3 (100%) 3 (100%) 6
ORSF cost 1 (3.3%) 0 (0%) 1
ORSF gbsg 1 (6.7%) 0 (0%) 1
ORSF hdfail 3 (100%) 3 (100%) 6
ORSF nafld1 1 (6.7%) 9 (60%) 10
ORSF uis 0 (0%) 1 (3.3%) 1
ORSF veteran 1 (3.3%) 0 (0%) 1
RRT dataFTR 5 (16.7%) 0 (0%) 5
RRT lung 8 (26.7%) 0 (0%) 8
RRT metabric 7 (46.7%) 0 (0%) 7
RRT nwtco 7 (46.7%) 0 (0%) 7
RRT ova 3 (10%) 0 (0%) 3
RRT tumor 3 (10%) 0 (0%) 3
MBSTCox child 3 (100%) 3 (100%) 6
MBSTCox dataSTR 0 (0%) 1 (3.3%) 1
MBSTCox hdfail 3 (100%) 3 (100%) 6
MBSTAFT hdfail 2 (66.7%) 0 (0%) 2
SSVM check_times 3 (100%) 0 (0%) 3
SSVM child 3 (100%) 0 (0%) 3
SSVM colrec 3 (100%) 0 (0%) 3
SSVM flchain 11 (73.3%) 0 (0%) 11
SSVM hdfail 3 (100%) 0 (0%) 3
SSVM nafld1 15 (100%) 0 (0%) 15
SSVM nwtco 8 (53.3%) 0 (0%) 8
SSVM ova 3 (10%) 0 (0%) 3
SSVM support 3 (100%) 0 (0%) 3
SSVM wa_churn 15 (100%) 0 (0%) 15
Appendix F Results per Dataset
@@
Figure 15: Per-dataset scores for learners tuned and evaluated with Harrell’s C
Figure 16: Per-dataset scores for learners tuned on Harrell’s C and evaluated with Uno’s C
Figure 17: Per-dataset scores for learners tuned and evaluated with ISBS
Figure 18: Per-dataset scores for learners tuned on ISBS evaluated with Harrell’s C
Appendix G Competing interests
No competing interest is declared.
Appendix H Data Availability Statement
The data underlying this article are available in the online supplement at https://github.com/slds-lmu/paper_2023_survival_benchmark.
Appendix I Author contributions statement
LB contributed to the planning and dataset selection, conducted the experiment and wrote the manuscript. JZ contributed to underlying software implementations and advised on model evaluation. RS conducted a pilot version of this study (Sonabend, 2021) which laid the foundation for this project and reviewed the manuscript. BB, AB, and MNW contributed to the planning, literature review, interpretation, and reviewed the manuscript.
Appendix J Acknowledgments
We are grateful for the package maintainers’ time supporting this effort regarding the model configurations used in this benchmark, in particular we thank Byron Jaeger, Matthias Schmid, Harald Binder, Cesaire Fouodo, Håvard Kvamme, Torsten Hothorn, Tong He, Chris Jackson, Terry Therneau, Trevor Hastie, and Jelle Goeman.